What fraction of the dogs are outside the camp, if ten dogs are outside the camp and four of the dogs are inside the camp

a. 4/10

b. 10/4

c. 6/7

d. 2 4/2


1
Expert's answer
2021-05-12T05:05:53-0400

Total number of dogs = 10 + 4 = 14

Fraction of the dogs are outside the camp "= \frac{10}{14} = \frac{5}{7}"